how to state the slope n y-intercept of the graph of this equation: y=-2/3-4
Answer by richwmiller(17219)

Are you missing something in your equation?
Should it be y=-(2/3)x-4 ?
BTW slope is m not n
If the equation is y=-(2/3)x-4 then the slope m = -2/3 and the y intercept =-4 or as an ordered pair (0,-4)